OpenStudy (anonymous):

Plot a table of values and then graph f(-x)={2+x, If -2=

OpenStudy (anonymous):

im going to redefine your table of values X=-x Y=2+x x can be between -2 to 0 X|Y 0| 1| 2|

OpenStudy (anonymous):

 thats the graph my text gives as an answer.

OpenStudy (anonymous):

are you use its f(-x)?

OpenStudy (anonymous):

f(x)=2+x x=x y=2+x x|y -2| -1| 0|

OpenStudy (anonymous):

the graph would be incorrect because you have a specific domain